# FADPNet: Frequency-Aware Dual-Path Network for Face Super-Resolution

### v1

@@ III-C High-Frequency Enhancement Block (HFEB)
In our method, HFEB specializes in high-frequency detail recovery. Specifically, for the high-frequency feature $\mathbf{F}_{H}\in\mathbb{R}^{C\times H\times W}$ obtained from frequency separation operator, we first apply a ${1\times 1}$ convolution to reduce channels to $\frac{C}{2}$ to minimizes computational costs. Then we use a ${1\times 1}$ convolution to expand channels to $C$ for follow-up operation. Subsequently, from a local enhancement perspective, features traverse a sequence of our High-Frequency Refinement (HFR), which progressively extracts hierarchical high-frequency patterns, strengthening details through iterative feature transformation. From a global modeling angle, interspersed among subsequent residual blocks, Depthwise Position-aware Attention (DPA) modules selectively enhance critical high-frequency facial attributes, while preserving facial realism through adaptive feature modulation. This synergistic operation establishes essential groundwork for achieving high-fidelity face super-resolution.
@@ III-C1 High-Frequency Refinement (HFR)
As shown in Fig. 4 (a), our HFR module operates in a recurrent manner, to iteratively refine high-frequency components—fine-grained facial details like eyes, eyebrows, and lips—over $r$ cycles. Given an input feature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, the module first extracts multi-scale spatial features through parallel depth-wise separable convolutions with complementary receptive fields:
@@
[…]
@@ III-C2 Depthwise Position-aware Attention (DPA)
[…]
@@
where $\mathrm{Conv}_{6},\mathrm{Conv}_{7}$ denote ${1\times 1}$ convolutions. However, we observe that using positional encoding alone did not yield the desired performance. To further refine the attention mechanism, we introduce a temperature generator module using a series of linear transformations to generate a temperature scale $\mathbf{temp}$, which $\mathbf{temp}$ can dynamically adjust attention scaling factors:
@@
[…]
@@ III-D Offsets mechanism
TABLE I: Ablation study of HFR on CelebA test [4].
Methods Params FLOPs PSNR SSIM
@@
TABLE III: Ablation study of ASSB is conducted on the Helen test set [46] by replacing it with three representative alternatives: self-attention (SA) [47], convolutional blocks (Conv), and Vision State Space Block (VSSB) [31].
@@
ASSB (Ours) 0.203M 0.125G 27.06 0.8064 0.4706 0.2167
@@
Figure 6: (a) Ablation studies of the effectiveness of frequency-specific modules on the CelebA [4] and Helen datasets [46]. “swap” is our model in which the low- and high-frequency blocks are swapped, resulting in a noticeable performance drop compared to our original model. (b) Feature map visualization and FSR results between our original model and swapped-frequency variant show our modules design for different frequency features produces sharp and accurate facial contours.
TABLE IV: Quantitative comparisons for $\times$8 SR on the CelebA and Helen test sets.
@@
FADPNet (Ours) 8.6M 39.6ms 28.21 0.8075 0.4876 0.1741 27.06 0.8064 0.4706 0.2167
TABLE V: Ablation study of our attention strategy on CelebA [4] and Helen [46] test set.
Methods CelebA Helen
PSNR /  SSIM PSNR /  SSIM
LFEB w/o SEB 28.16 /  0.8067 26.99 /  0.8043
HFEB w/o DPA 28.04 /  0.8032 26.80 /  0.7981
w/o offsets 28.17 /  0.8073 27.00 /  0.8050
Ours 28.21 /  0.8075 27.06 /  0.8064
Figure 7: Visual comparisons for $\times$8 SR on the CelebA test set [4]. Our FADPNet can reconstruct clear face images.
IV Experiments
IV-A Datasets and Evaluation Metrics
We employ the CelebA [4] dataset for training and evaluate our model on three benchmarks: CelebA [4], Helen [46], and SCface [50]. Therefore, we crop the image based on its center point and resize it to ${128\times 128}$ pixels, serving as the high-resolution (HR) ground truth. Corresponding low-resolution (LR) inputs (${16\times 16}$ pixels) are generated by downsampling HR images via bicubic interpolation. The training set comprises 18,000 images from CelebA, with 1,000 and 200 images reserved for testing and validation, respectively. For cross-dataset evaluation, we test on 50 samples from Helen and the SCface dataset without fine-tuning. We assess reconstruction quality using four widely adopted metrics: PSNR and SSIM [51] for pixel-level fidelity, LPIPS [52] for perceptual similarity, and VIF [53] for texture preservation.
IV-B Implementation Details
Our FADPNet is implemented in PyTorch and trained on an $NVIDIA\quad RTX\quad 3090\quad$ GPU. We initialize the model with 32 channels and design its architecture across three distinct resolution levels. Specifically, the number of low-frequency modules is 2 at each level. Meanwhile, the number of HEFB modules is configured as follows: 2 in $\mathbb{R}^{\mathrm{C}\times\mathrm{H}\times\mathrm{W}}$ stage, 3 in $\mathbb{R}^{2\mathrm{C}\times\frac{\mathrm{H}}{2}\times\frac{\mathrm{W}}{2}}$ stage, and 4 in $\mathbb{R}^{4\mathrm{C}\times\frac{\mathrm{H}}{4}\times\frac{\mathrm{W}}{4}}$ stage. Optimization is performed using the Adam optimizer (${\beta_{1}=0.9}$ and ${\beta_{2}=0.99}$) with an initial learning rate of ${2\times 10^{-4}}$ The loss function combines reconstruction loss, and training converges within 150 epochs using a batch size of 16.
Figure 8: Visual comparisons for $\times$8 SR on the Helen test set [46]. Our FADPNet can reconstruct clear face images.
@@ IV-B Implementation Details
IV-C Ablation Studies
IV-C1 Effectiveness of HFR
IV-C2 Effectiveness of DPA
To comprehensively evaluate the individual and synergistic effects of components in our proposed Depthwise Position-aware Attention (DPA) module, we perform systematic ablation studies on the Helen test set [1], with detailed results presented in Table II. Our investigation focuses on three critical elements: the learnable temperature generator (T_gen), fixed temperature parameter (T_para), and positional embedding (P_emb), analyzing their impacts on both reconstruction quality and model complexity.
IV-C3 Effectiveness of ASSB
To evaluate the suitability of the Attentive State Space Block (ASSB) for low-frequency modeling in our method, we conduct ablation studies by replacing it with three representative alternatives: self-attention (SA) [47], residual block (CNN) [54], and Vision State Space Block (VSSB) [31]. As shown in Table III, our full model with ASSB achieves the best overall performance, indicating superior perceptual quality and structural fidelity.
It is worth noting that ASSB is not our original design, but a recently proposed state-space module tailored for long-range dependency modeling. In our framework, we integrate ASSB within LFEB to exploit its strength in global structure preservation. Unlike Conv, which tends to focus on local or high-frequency features, ASSB’s non-causal state-space formulation excels at capturing smooth, spatially consistent facial regions such as contours and skin tones. It’s built-in low-rank prompt decomposition also aids in suppressing high-frequency noise and promotes stable propagation across homogeneous areas. Compared to the more general-purpose VSSB, ASSB offers a lighter structure and better frequency selectivity when dealing with low-frequency signals. Despite using fewer parameters (0.203M vs. 0.269M), ASSB outperforms or matches VSSB in all metrics, demonstrating its superior efficiency-to-performance ratio when applied to low-frequency pathways. These results collectively justify our choice of ASSB as the most effective backbone for global low-frequency modeling in our frequency-aware architecture. Its integration helps maintain large-scale facial coherence and mitigates structural misalignment—challenges that conventional transformer or Conv-based designs struggle to address.
IV-C4 Study of the architecture of low- and high-frequency modules
To investigate the effectiveness of low- and high-frequency architecture, we conduct an ablation study on the CelebA and Helen datasets by interchanging the positions of the low-frequency enhancement block (LFEB) and high-frequency enhancement block (HFEB). As illustrated in Fig. 6(a), this modification leads to a consistent drop in PSNR across both datasets, highlighting the critical importance of the original frequency-aware design. The original model achieves superior reconstruction performance by appropriately allocating low- and high-frequency processing to their respective branches. In contrast, the swapped configuration disrupts this balance, resulting in suboptimal representation and degraded image quality. Fig. 6(b) visually compares the original and the swapped variants. The results clearly show that the proposed model preserves facial structures and fine details more effectively, yielding sharper contours and more realistic textures. Conversely, the swapped version produces noticeably blurrier outputs, especially around facial components like eyes and mouth, further validating the necessity of maintaining the original design of frequency-specific modules.
@@ IV-C5 Study of local-global attention strategy
Figure 10: Visual comparison on SCface [50] of real surveillance scenarios for $\times$8 SR. Our FADPNet can reconstruct clear face components.
TABLE VI: Quantitative comparison on face average similarity with existing methods on SCface [50] of real surveillance scenarios.
@@ IV-D Comparison with Other Methods
In this section, we conduct a comprehensive comparison of our proposed method against state-of-the-art (SOTA) techniques in FSR, encompassing general image SR methods like RCAN [54], Restormer-M [47] and FreqFormer [36], specialized FSR approaches such as FSRNet [13], DIC [14], SPARNet [16], AD-GNN [19], SISN [18], LAAT [20], SFMNet [22], ELSFace [48], and SPADNet [49]. All methods are trained on the CelebA [4] dataset under identical settings—including preprocessing, loss functions, and optimizer configurations—to eliminate implementation biases and ensure direct attribution of performance differences to architectural innovations.
@@ IV-D1 Comparison on the CelebA dataset and the Helen dataset
We present the quantitative results for the CelebA and Helen test datasets in Table IV. The best and second-best results are highlighted in bold and underlined, respectively. These results demonstrate the strength of our architecture in balancing structural accuracy and perceptual quality. To further evaluate the generalization ability of our method, we directly apply the model trained on CelebA to the Helen test set. The quantitative results for $\times$8 FSR are also presented in Table IV, showing that our method comprehensively outperforms other approaches. Despite having only 8.6M parameters (similar to SFMNet [22]), ur FADPNet achieves 39.6ms inference speed, which is 1.8× faster than FreqFormer [36] (75ms) and comparable to lightweight methods like SPARNet [16] (36.6ms). Compared to ELSFace [48], despite the similar parameter counts (8.6M vs. 8.7M), FADPNet improves PSNR by 0.25dB on CelebA and 0.36dB on Helen. As summarized in Fig. 7 and Fig. 8, while competing methods often produce blurred facial contours and distorted features (e.g., ill-defined eyes), our reconstructions successfully preserve sharp edges, facial symmetry, and fine-grained details (e.g., eyeball), closely aligning with the ground-truth HR images. As shown in Fig. 9, we also use the face parsing model [56] to analyze the FSR results of different models on the Helen dataset. The visual results clearly demonstrate that our method produces the most complete and accurate facial structure, preserving better face details than competing models. In summary, while existing methods struggle with blurred shapes and significant detail loss, our model effectively preserves facial contours and fine details. This highlights not only the robustness of our approach across different datasets but also its superiority over SOTA techniques in both performance and efficiency.
IV-D2 Comparison on real-world surveillance scenarios
Restoring facial details from real-world surveillance imagery remains a formidable challenge due to inherent low resolution, uncontrolled lighting, and sensor noise—factors that are often inadequately captured in synthetic benchmarks. To address this, we conduct experiments using low-quality face images from the SCface dataset [50], which inherently represents real-world surveillance scenarios without manual downsampling. As illustrated in Fig. 10, visual comparisons reveal that existing methods often struggle to reconstruct critical facial features due to the challenges posed by noisy, low-resolution inputs, leading to blurred contours and loss of texture details. In contrast, our FADPNet leverages adaptive feature separation and dual frequency enhancement, effectively restoring sharper facial structures, finer textures, and more natural facial symmetry. Beyond visual fidelity, we assess the impact on downstream tasks such as face matching, comparing the similarity between restored faces and their corresponding high-definition reference images. Table VI presents the quantitative results, demonstrating that our approach consistently achieves higher similarity scores across multiple test cases. This confirms its practical applicability in challenging surveillance scenarios.

### v2

@@ III-C High-Frequency Enhancement Block (HFEB)
In our method, the High-Frequency Enhancement Block (HFEB) targets high-frequency detail recovery. Given high-frequency features $\mathbf{F}_{H}\in\mathbb{R}^{C\times H\times W}$ from the frequency separation operator, we first apply a $1\times 1$} convolution to reduce channels to $\frac{C}{2}$, then expand them back to $C$ to reduce computation. Locally, features pass through the High-Frequency Refinement (HFR) module, which progressively extracts hierarchical high-frequency patterns and strengthens details. Globally, Depthwise Position-aware Attention (DPA) modules interleaved with residual blocks selectively enhance critical high-frequency attributes while preserving realism. This combination forms the basis for high-fidelity FSR.
@@ III-C1 High-Frequency Refinement (HFR)
As shown in Fig. 4 (a), our HFR module operates recurrently, to iteratively refine high-frequency components—fine-grained facial details like eyes, eyebrows, and lips—over $r$ cycles. Given an input feature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, the module first extracts multi-scale spatial features through parallel depth-wise separable convolutions with complementary receptive fields:
@@
$\mathbf{X}_{\text{3}}=\mathrm{Conv}_{\text{4}}\cdot\mathit{CS}(\mathrm{Conv}_{\text{3}}\cdot\mathbf{X}_{\text{cat}}),$ (11)
where $\mathrm{Conv}_{\text{3}}$ is a $1\times 1$} convolution for channel expansion, $\mathit{CS}(\cdot)$ denotes channel-shuffle and $\mathrm{Conv}_{\text{4}}$ is a $1\times 1$} convolution for channel reduction. Combined with channel shuffle, this operation dynamically permutes channel subgroups to enhance inter-channel communication.
@@ III-C2 Depthwise Position-aware Attention (DPA)
$\mathit{Q,K,V}=\mathit{RS}\left(\mathrm{DWConv}_{3\times 3}\left(\mathrm{Conv}_{\text{5}}\left(\mathbf{Y}_{1}\right)\right)\right),$} (12)
where $\mathrm{Conv}_{\text{5}}$ is a $1\times 1$} convolution that expands channels from $C$ to $3C$, $\mathrm{DWConv}_{3\times 3}$ denotes a $3\times 3$} depthwise convolution and $\mathit{RS}$ denotes a reshape and split operator. To enhance positional modeling capabilities, we introduce a dynamical positional encoding mechanism. The input feature map is processed through a sub-network consisting of two $1\times 1\times 1$ convolutions with $\operatorname{GELU}$ activation in between:
@@
where $\mathrm{Conv}_{6},\mathrm{Conv}_{7}$ denote $1\times 1$} convolutions. However, we observe that using Positional Embedding alone did not yield the desired performance. To further refine the attention mechanism, we introduce a Temperature Generator module using a series of linear transformations to generate a temperature scale $\mathbf{temp}$, which can dynamically adjust attention scaling factors:
@@
$\operatorname{Attn}\mathit{\{Q,K,V\}}=\mathit{V}\cdot\operatorname{RELU}(\mathit{QK}^{T}\cdot\mathbf{temp})+\sigma(\mathbf{pos}),$ (15)
@@ III-D Offsets Mechanism
III-E Loss Function
Given a training pair $\left\{\boldsymbol{I}_{LR},\boldsymbol{I}_{HR}\right\}$, our model predicts the super-resolved image $\boldsymbol{I}_{SR}$ from the low-resolution input $\boldsymbol{I}_{LR}$. The overall training objective is defined as
$\mathcal{L}_{total}=\mathcal{L}_{pixel}=\left\|\boldsymbol{I}_{SR}-\boldsymbol{I}_{HR}\right\|_{1},$ (17)
where $\lambda=1$, and $\boldsymbol{I}_{HR}$ denotes the corresponding high-resolution ground truth. No additional perceptual, adversarial, or frequency-domain loss is employed in training.
IV Experiments
IV-A Datasets and Evaluation Metrics
We employ the CelebA [3] dataset for training and evaluate our model on three benchmarks: CelebA [3], Helen [41], and SCface [42]. Therefore, we crop the image based on its center point and resize it to $128\times 128$} pixels, serving as the high-resolution (HR) ground truth. Corresponding low-resolution (LR) inputs ($16\times 16$} pixels) are generated by downsampling HR images via bicubic interpolation. The training set comprises 18,000 images from CelebA, with 1,000 and 200 images reserved for testing and validation, respectively. For cross-dataset evaluation, we test on 50 samples from Helen and the SCface dataset without fine-tuning. We assess reconstruction quality using four widely adopted metrics: PSNR and SSIM [43] for pixel-level fidelity, LPIPS [44] for perceptual similarity, and VIF [45] for texture preservation. For runtime evaluation, the reported runtime is computed as the average inference time over 1,000 test images with batch size 1. Inference is performed in FP32 precision without warmup runs, and the reported timing only includes network forward inference, excluding preprocessing and data transfer.
IV-B Implementation Details
Our FADPNet is implemented in PyTorch and trained on an $NVIDIA\quad RTX\quad 3090\quad$ GPU. We initialize the model with 32 channels and design its architecture across three distinct resolution levels. Specifically, the number of low-frequency modules is 2 at each level. Meanwhile, the number of HFEB modules is configured as follows: 2 in $\mathbb{R}^{\mathrm{C}\times\mathrm{H}\times\mathrm{W}}$ stage, 3 in $\mathbb{R}^{2\mathrm{C}\times\frac{\mathrm{H}}{2}\times\frac{\mathrm{W}}{2}}$ stage, and 4 in $\mathbb{R}^{4\mathrm{C}\times\frac{\mathrm{H}}{4}\times\frac{\mathrm{W}}{4}}$ stage. Optimization is performed using the Adam optimizer ($\beta_{1}=0.9$} and $\beta_{2}=0.99$}) with an initial learning rate of $2\times 10^{-4}$}. The model is optimized using the reconstruction loss defined in Section III-E, and training converges within 150 epochs using a batch size of 16. During training, standard data augmentation is applied, including random horizontal flipping and random scaling with a factor sampled from $[1.0,1.3]$. The low-resolution input is generated by bicubic downsampling of the HR image, followed by a bicubic upsampling to the target spatial size.
IV-C Ablation Studies
IV-C1 Effectiveness of HFR
IV-C2 Effectiveness of DPA
To comprehensively evaluate the individual and synergistic effects of components in our proposed Depthwise Position-aware Attention (DPA) module, we perform systematic ablation studies on the Helen test set, with detailed results presented in Table II. Our investigation focuses on three critical elements: the learnable Temperature Generator (T_gen), fixed Temperature Parameter (T_para), and Positional Embedding (P_emb), analyzing their impacts on both reconstruction quality and model complexity.
TABLE I: Ablation study of HFR on CelebA test set [3].
[…]
@@ IV-C2 Effectiveness of DPA
TABLE III: Ablation study of ASSB on the Helen test set [41] by replacing it with three representative alternatives: Self-Attention (SA) [46], Convolutional block (CNN), and Vision State Space Block (VSSB) [8].
@@
ASSB 0.203M 0.125G 27.06 0.8064 0.4706 0.2167
TABLE IV: Component-level ablation of ASSB on CelebA test set [3] and Helen test set [41].
Methods CelebA Helen
$uparrow PSNR\uparrow / SSIM\uparrow$
w/o prompt pool 28.16 / 0.8061 26.99 / 0.8035
w/o routing 28.18 / 0.8069 27.00 / 0.8040
w/o semantic reorder 28.18 / 0.8074 26.99 / 0.8039
w/o low-rank prompt 28.19 / 0.8074 27.02 / 0.8045
Full ASSB 28.21 / 0.8075 27.06 / 0.8064
IV-C3 Effectiveness of ASSB
To evaluate the suitability of the Attentive State Space Block (ASSB) for low-frequency modeling in our method, we conduct ablation studies by replacing it with three representative alternatives: Self-Attention (SA) [46], Convolutional block (CNN), and Vision State Space Block (VSSB) [8]. As shown in Table III, our full model with ASSB achieves the best overall performance, indicating superior perceptual quality and structural fidelity.
It is worth noting that ASSB is not our original design, but a recently proposed state-space module tailored for long-range dependency modeling. In our framework, we integrate ASSB within LFEB to exploit its strength in global structure preservation. Unlike CNN, which tends to focus on local or high-frequency features, ASSB’s non-causal state-space formulation excels at capturing smooth, spatially consistent facial regions such as contours and skin tones. Its built-in low-rank prompt decomposition also aids in suppressing high-frequency noise and promotes stable propagation across homogeneous areas. Compared to the more general-purpose VSSB, ASSB offers a lighter structure and better frequency selectivity when dealing with low-frequency signals. Despite using fewer parameters (0.203M vs. 0.269M), ASSB outperforms or matches VSSB in all metrics, demonstrating its superior efficiency-to-performance ratio when applied to low-frequency pathways. These results collectively justify our choice of ASSB as the most effective backbone for global low-frequency modeling in our frequency-aware architecture. Its integration helps maintain large-scale facial coherence and mitigates structural misalignment—challenges that conventional transformer or Conv-based designs struggle to address.
We also conduct ablation studies by removing or simplifying individual components in ASSB while keeping the rest of the architecture unchanged. These components include the prompt pool, routing strategy, semantic reorder, and low-rank prompt design. Together, these ablations cover the key prompt-guided and sequence-organization elements in ASSB, where the semantic reorder variant also reflects the contribution of the non-causal scanning behavior introduced by sequence reorganization. As shown in Table IV, simplifying any of these components leads to a performance drop on both datasets. This indicates that the improvement brought by ASSB does not arise from naive module stacking, but from the complementary effects of its internal components. In particular, the prompt-related design, routing strategy, semantic reorder, and low-rank prompt adaptation all contribute to the final performance. Meanwhile, the number of parameters and FLOPs remains nearly unchanged across variants, since these ablations mainly affect internal feature organization and prompt generation strategies rather than the overall backbone width or depth.
TABLE V: Ablation study of our attention strategy on CelebA test set [3] and Helen test set [41].
Methods CelebA Helen
$uparrow PSNR\uparrow / SSIM\uparrow$
LFEB w/o SEB 28.16 / 0.8067 26.99 / 0.8043
HFEB w/o DPA 28.04 / 0.8032 26.80 / 0.7981
w/o offsets 28.17 / 0.8073 27.00 / 0.8050
Ours 28.21 / 0.8075 27.06 / 0.8064
TABLE VI: Ablation study of the offset-based alignment mechanism on CelebA test set [3] and Helen test set [41].
Meehods CelebA Helen
$uparrow PSNR\uparrow / SSIM\uparrow$
w/o offset 28.17 / 0.8073 27.00 / 0.8050
w/ fixed warp 28.20 / 0.8068 27.01 / 0.8047
w/ zero-offset warp 28.18 / 0.8069 27.00 / 0.8033
w/ conv alignment 28.18 / 0.8065 27.00 / 0.8049
Offset (Ours) 28.21 / 0.8075 27.06 / 0.8064
@@ IV-C3 Effectiveness of ASSB
Figure 6: (a) Ablation studies of frequency-specific modules on the CelebA [3] and Helen datasets [41]. “swap” is our model in which the low- and high-frequency blocks are swapped, resulting in a noticeable performance drop compared to our original model. (b) Feature map visualization and FSR results between our original model and swapped-frequency variant show that our module design for different frequency features produces sharp and accurate facial contours.
IV-C4 Study of the architecture of low- and high-frequency modules
To investigate the effectiveness of low- and high-frequency architecture, we conduct an ablation study on the CelebA and Helen datasets by interchanging the positions of LFEB and HFEB. As illustrated in Fig. 6(a), this modification leads to a consistent drop in PSNR across both datasets, highlighting the critical importance of the original frequency-aware design. The original model achieves superior reconstruction performance by appropriately allocating low- and high-frequency processing to their respective branches. In contrast, the swapped configuration disrupts this balance, resulting in suboptimal representation and degraded image quality. Fig. 6(b) visually compares the original and the swapped variants. The results clearly show that the proposed model preserves facial structures and fine details more effectively, yielding sharper contours and more realistic textures. Conversely, the swapped version produces noticeably blurrier outputs, especially around facial components like eyes and mouth, further validating the necessity of maintaining the original design of frequency-specific modules.
TABLE VII: Quantitative comparisons for $\times$8 FSR on CelebA test set [3] and Helen test set [41].
@@ IV-C4 Study of the architecture of low- and high-frequency modules
FADPNet-S 4.8M 30.0ms 28.13 0.8055 0.4848 0.1807 26.95 0.8015 0.4652 0.2233
FADPNet 8.6M 39.6ms 28.21 0.8075 0.4876 0.1741 27.06 0.8064 0.4706 0.2167
Figure 7: Visual comparisons for $\times$8 FSR on CelebA test set [3] and Helen test set [41]. Please zoom in for more details.
@@ IV-C5 Study of local-global attention strategy
To provide a more informative analysis of the offset mechanism, we further compare the proposed offset-based alignment module with several alternatives, including: (1) removing the alignment branch entirely, (2) replacing the learned offsets with a fixed warp, (3) using zero-offset warping, and (4) replacing it with a lightweight convolutional alignment branch. As shown in Table VI, replacing the learned offset warp with a fixed warp, zero-offset warp, or convolution-based alignment fails to consistently mitigate cross-scale feature misalignment and leads to inferior overall performance across datasets and metrics. This indicates that the improvement does not simply come from adding an extra refinement branch or a generic warping operator, but specifically from the learned offset-based alignment itself.
Figure 9: Visual comparison on SCface [42] of real surveillance scenarios for $\times$8 SR. Our FADPNet achieves better FSR.
TABLE VIII: Comparisons on face average similarity with existing methods on SCface [42] of real surveillance scenarios.
@@ IV-D Comparison with Other Methods
We compare our method with existing FSR methods, encompassing general image SR methods like Restormer-M [46], FreqFormer [33] and FreMamba [49], specialized FSR approaches such as FSRNet [13], DIC [14], SPARNet [16], AD-GNN [19], SISN [18], LAAT [20], SFMNet [22], ELSFace [47], SPADNet [48] and DMNet [34]. All methods are trained on the CelebA [3] dataset under identical settings—including preprocessing, loss functions, and optimizer configurations—to eliminate implementation biases.
@@ IV-D1 Comparison on the CelebA dataset and the Helen dataset
Table VII reports quantitative results on CelebA and Helen, with the best and second-best scores marked in bold and underlined, respectively. Our model, trained on CelebA and directly applied to Helen, achieves state-of-the-art $\times$8 FSR performance, consistently balancing structural accuracy and perceptual quality. With only 8.6M parameters—similar to SFMNet [22]—FADPNet runs at 39.6 ms, 1.8× faster than FreqFormer [33] (75 ms) and comparable to SPARNet [16] (36.6 ms). Compared to ELSFace [47], it improves PSNR by 0.25dB on CelebA and 0.36dB on Helen. Compared with those recent frequency-aware methods such as FreMamba [49] and DMNet [34], our FADPNet still maintains the best performance and faster inference speed on both CelebA and Helen datasets, further demonstrating the effectiveness of the proposed frequency-dispatched design. Our lighter variant, denoted as FADPNet-S (the channel width is reduced, while the overall frequency-aware dual-path framework is preserved), still achieves competitive performance with lower computational cost. As also shown in Fig. 7, competing methods often yield blurred contours or distorted features (e.g., ill-defined eyes), whereas our reconstructions preserve sharp edges, facial symmetry, and fine details (e.g., eyeball), closely aligning with the ground-truth HR images. Face parsing results on Helen (as shown in Fig. 8) further confirm that FADPNet produces the most complete and accurate facial structures. Overall, our method outperforms prior methods in both performance (PSNR) and efficiency (Params-FLOPs-Speed) while maintaining robust generalization across datasets.
IV-D2 Comparison on real-world surveillance
Restoring facial details from real surveillance imagery is challenging due to LR, uncontrolled lighting, and sensor noise—conditions often underrepresented in synthetic benchmarks. We evaluate on SCface [42], which contains naturally degraded surveillance images without manual downsampling. As shown in Fig. 9, existing methods frequently produce blurred contours and lose texture details under such conditions. In contrast, FADPNet, with adaptive feature separation and dual-frequency enhancement, recovers sharper structures, finer textures, and natural facial symmetry. We further assess downstream performance in face matching by comparing restored faces with high-definition references. Table VIII shows that FADPNet consistently achieves higher similarity scores, confirming its practical effectiveness in challenging surveillance scenarios.